OpenStudy (anonymous):

Find the y-intercept for this equation: 3x - 5y = 30

OpenStudy (photon336):

Isolate Y by subtracting 3x to get the following \[5y = 3x-30 \] then isolate y by dividing both sides by 5 \[y = \frac{ 3 }{ 5 }x - 6\] this is the equation of a line which is below. \[y = m(x)+b \] m = slope change in y/x and b is the y intercept. \[y = \frac{ 3 }{ 5 }(0) - 6 = -6 \] how do we find the y intercept? we do this by plugging x = 0 into our equation. that's the point where the graph touches the y axes and x = 0 who when x = 0 y = -6

OpenStudy (photon336):

@siahniya16

OpenStudy (anonymous):

@Photon336 Tahnks